ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 81R50,81Q60,81Q05,81V19 Ndimubandi, Jean, Université du Burundi The q-Dirac oscillator: the context of radial potentials. The Dirac equation is discussed in connection with the theory of quantum deformations. We point out the so-called q-Dirac oscillator and we extend our considerations to radial potentials.